OpenStudy (anonymous):

f(x)=2x=2 and g(x)=x^2-1 find in simplest form f(-2) g(k+1)

OpenStudy (anonymous):

I just want to make sure that you meant \(\large f(x) = 2x\color{red}{+}2\). :-)

OpenStudy (anonymous):

yeah sorry

OpenStudy (anonymous):

Note that if \(\large f(x)=2x+2\), then \(\large f(-2) = 2(-2)+2 = \ldots \). Can you take things from here? Note that if \(\large g(x) = x^2-1\), then \(g(k+1) = (k+1)^2-1 = (k+1)(k+1)-1\). All you need to do now to finish this part is FOIL out the product there and simplify. Can you take things from here? I hope this makes sense! :-)

OpenStudy (anonymous):

okay I understand! Thankyou!